Density of a gas at 276k and 380mm pressure is 1.5grams/litre. Weight of 1.25 mole of the gas in grams is

Grade:12th pass

1 Answers

aditya kulkarni
6 years ago
using density=pm/rt where p=pressure , m= molar mass, r=universal gas constant, t= temperature, we get-
1.5=(1.5*m)/(0.0821*276)
on solving this eqn, we get     m=70
1 mole = 70gram
therefore 1.25 mole= 1.25*70=87.5gram